OpenStudy (anonymous):

Derivative of the absolute value of x-1?

OpenStudy (anonymous):

what is the derivative of \(-x+1\)?

OpenStudy (anonymous):

-1

OpenStudy (raden):

derive implicitly y = |x-1| y^2 = (x-1)^2 2y y' = 2(x-1) y' = 2(x-1)/(2y) = (x-1)/|x-1|

OpenStudy (anonymous):

and what is the derivative of \(x-1\) ?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

\[f(x) = |x-1| = \left\{\begin{array}{rcc} -x+1& \text{if} & x <1 \\ x-1& \text{if} & x \geq 14\end{array} \right. \]

OpenStudy (anonymous):

typo there but you get the idea right?

OpenStudy (anonymous):

I think so. from there how would you get the value of x on the clos interval 0 to 4?

OpenStudy (anonymous):

\[f(x) = |x-1| = \left\{\begin{array}{rcc} -x+1& \text{if} & x <1 \\ x-1& \text{if} & x \geq 1\end{array} \right.\] \[f(x) = |x-1| = \left\{\begin{array}{rcc} -1& \text{if} & x <1 \\ 1& \text{if} & x \geq 1\end{array} \right.\]

OpenStudy (anonymous):

that should have been \[f'(x) = |x-1| = \left\{\begin{array}{rcc} -1& \text{if} & x <1 \\ 1& \text{if} & x > 1\end{array} \right.\]

OpenStudy (anonymous):

also \(f'(1)\) does not exist as \(1\neq -1\)

OpenStudy (anonymous):

i am not sure what you mean by "the value of x on the closed interval"